Calculate the mass of iodine that must be

reacted with excess phosphorus to produce

5.00 g of phosphorus(III) iodide according to

the equation below.

2P+3I2 → 2PI3


1
Expert's answer
2021-07-20T12:10:01-0400

Solution.

"xg" "500g"

"2P+3I_2 \u2192 2PI_3"

"3 Mr(I_2)" "2Mr(PI_3)"

"\dfrac{x}{3Mr(I_2)}=\dfrac{500}{2Mr(PI_3)};"

"x=\dfrac{500\sdot3Mr(I_2)}{2Mr(PI_3)};"

"x=\dfrac{500\sdot3\sdot126.9\sdot2}{2\sdot(31+126.9\sdot3)}=462g;"

Answer: "462g."